would really appreciate help on this one!
solve for x:
2log[base 4]x-log[base 4](x-1)=1

Answer by stanbon(75887)

2log[base 4]x-log[base 4](x-1)=1
log(4)[x^2/(x-1)=1
x^2/(x-1)=4
x^2-4x+4=0
(x-2)^2=0
x=2 with multiplicity 2
Cheers,
Stan H.